randy invested his inheritance in an account that paid 6.9% interest, coupounded continuously. after 5 years, he found that he now had $52,565.56. what was the orginal amount of his inheritance?

Answer by lwsshak3(11628)

randy invested his inheritance in an account that paid 6.9% interest, coupounded continuously. after 5 years, he found that he now had $52,565.56. what was the orginal amount of his inheritance?
**
Formula for continuous compounding:A= Pe^rt, P=initial investment, r=rate of interest, A=amt after t years
52565.56=Pe^(.069*5)=Pe^.345
P=52565.56/e^.345=37228
orginal amount of his inheritance=$37,228.00..